OpenStudy (anonymous):

he length of a rectangle is 16 cm longer than twice its width. The perimeter is 122 cm. Find the dimensions of the rectangle.

OpenStudy (anonymous):

2(2x+16 + x) = 122 3x + 16 = 61 3x = 45 x = 15 and 2x+16 = 46 length = 46 ft and width = 15 ft <==ANSWER

OpenStudy (anonymous):

Welcome to Open Study! You can give medals by clicking 'Best Response'